OpenStudy (anonymous):

Ln(1/x - 2 ) - ln x =1

OpenStudy (anonymous):

raise everything to

OpenStudy (anonymous):

e

OpenStudy (anonymous):

1/x-2 - x = e

OpenStudy (anonymous):

Then after that ?

OpenStudy (anonymous):

solve for x

OpenStudy (anonymous):

1-x+2/x(x+2)=e

OpenStudy (anonymous):

1+x/x(x+2)=e

OpenStudy (anonymous):

Then?